the amount of heat required to raise the temperature of 100gm water from 20 degree celsius to 40 deghree celsius will be
 

Grade:12

2 Answers

Vikas TU
4 years ago
We use the heat equation. Q=mcΔT.
Where, m is the mass of water in kg, c is specific heat capacity in J∙kg−1°C−1 and ΔT the temperature difference.
In this case ,
Q=0.1×4180×(40–20)
= 8360 J, the amount of heat required to do that.
Khimraj
4 years ago
𝑄=𝑚𝑐Δ𝑇 Where, m is the mass of water in kg, c is specific heat capacity in J•kg−1°C−1 and ΔT the temperature difference. Q=0.1×4180×(40–20) = 8360 J